If P(B)=3/5 ,P(A/B)=1/2 and P(AUB)=4/5 , then P(AUB)’+P(A’​​^B)=
(i)1/5
(ii)4/5
(iii)1/2
(iv)11/10

Grade:12

1 Answers

Arun
4 years ago
Dear student
 
P(A/B) = ½
P(A ^ B )/ P(B) = ½
Hence P (A^B) = 3/10
 
Now
 
P (A U B)' + P (A' ^B)
 = 1 - P(AUB) + P (B) - P (B ^ A)
 = 1 - 4/5 + 3/5 - 3/10
 = (10 - 8 + 6 -3)/10 = 5/10 = ½
Hence option C is correct